OpenStudy (anonymous):

David opens a bank account with $140 and then deposits $150 each month. Louis opens a bank account with $260 and deposits $120 each month. How many months will it take for David and Louis to have equal amounts of money in their accounts? A. 2 B. 4 C. 5 D. 7

OpenStudy (anonymous):

@TheFlash17 plz help me

OpenStudy (anonymous):

Its 4

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

x= number of months that it takes for David and Louis to have equal amounts of money. Then: 140+x*150=260+x*120 \\ 150x-120x=260-140 \\ 30x=120 \\ \boxed{\bold{x=4}} Answer: it takes 4 months --> OPTION B

OpenStudy (anonymous):

ohhh... okay thank you so much

OpenStudy (anonymous):

ur quick

OpenStudy (texaschic101):

140 + 150m = 260 + 120m 150m - 120m = 260 - 140 30m = 120 m = 120/30 m = 4

OpenStudy (anonymous):

okay can someone plz medal @texaschic101

OpenStudy (texaschic101):

thats not necessary...but thanks anyways

OpenStudy (anonymous):

@texaschic101 there you go

OpenStudy (texaschic101):

thx :)

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

thx for yalls help